# Matrix Completion with Convex Optimization and Column Subset Selection

### v3

@@ Abstract
We introduce a two-step method for the matrix recovery problem. Our approach combines the theoretical foundations of the Column Subset Selection and Low-rank Matrix Completion problems. The proposed method, in each step, solves a convex optimization task. We present two algorithms that implement our Columns Selected Matrix Completion (CSMC) method, each dedicated to a different size problem. We performed a formal analysis of the presented method, in which we formulated the necessary assumptions and the probability of finding a correct solution. In the second part of the paper, we present the results of the experimental work. Numerical experiments verified the correctness and performance of the algorithms. To study the influence of the matrix size, rank, and the proportion of missing elements on the quality of the solution and the computation time, we performed experiments on synthetic data. The presented method was applied to two real-life problems problems: prediction of movie rates in a recommendation system and image inpainting. Our thorough analysis shows that CSMC provides solutions of comparable quality to matrix completion algorithms, which are based on convex optimization. However, CSMC offers notable savings in terms of runtime.
Keywords Matrix completion  $\cdot$ Column Subset Selection  $\cdot$ Low-rank models  $\cdot$ Image inpainting  $\cdot$ Recommendation system  $\cdot$ Nuclear Norm  $\cdot$ Convex Optimization
@@ 1 Introduction
Images, time series, graphs, and recommendation systems are just a few examples of data types with natural representation as real-valued matrices. In practice, these matrices often miss entries due to data collection costs, sensor errors, or privacy considerations. To fully exploit the potential of matrix-based datasets, we need to address the missing data problem.
Matrix completion (MC) has found applications in diverse fields such as signal processing, sensor localization, medicine and genomics [1, 2, 3, 4]. Its significance was highlighted in 2006 with the Netflix Prize challenge [5], aimed at improving recommendation systems by predicting unknown ratings. This task involves completing the sparse user-movie rating matrix $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$.
Singular Value Decomposition (SVD) is commonly employed to exploit the high correlation in the data,
@@
where $\mathbf{U}\in\mathbb{R}^{n_{1}\times n_{1}}$, $\mathbf{V}\in\mathbb{R}^{n_{2}\times n_{2}}$ are orthogonal, and $\mathbf{\Sigma}\in\mathbb{R}^{n_{1}\times n_{2}}$ is non-negative diagonal matrix of singular values, ordered according to their magnitude. Singular values capture the variance in the data. The rank of a matrix is determined by the count of its non-zero singular values. A common approach is to find the matrix with the lowest rank that matches the observed entries. However, directly minimizing the rank is NP-hard [6, 5]. Instead, minimizing the nuclear norm, which is the sum of the singular values of the matrix, is often pursued [7, 8, 9]. Let $\Omega\subseteq\{1,\cdots,n_{1}\}\times\{1,\cdots n_{2}\}$ be the set of indices of the known entries, the convex optimization problem for the exact matrix completion has the form,
@@ 2.3 Our contributions
    We introduce a two-staged matrix completion method, CSMC, dedicated to recovering rectangular matrices with one direction significantly larger than the other ($n_{1}<<n_{2}$).
@@
    We present results of several numerical experiments on synthetic and real-world datasets to verify theoretical results. The performance of the presented algorithms is compared with commonly used matrix completion algorithms described in the literature. Additionally, we apply the CSMC method to two real-life scenarios: movie recommendation and image inpainting.
@@ Results
Bridge restored with CSNN-0.5
Figure 8: Image inpainting; NN and CSNN-0.5 algorithms, $\rho=20\%$ known entries.
@@ 5 Conclusion
This paper introduces the Columns Selected Matrix Completion (CSMC) method, devised to enhance the time efficiency of nuclear norm minimization algorithms for low-rank matrix completion tasks. Through a two-staged approach, CSMC strategically applies a low-rank matrix completion algorithm to a reduced problem size in the first stage, followed by the minimization of least squares error in the second stage. Theoretical analysis provides provable guarantees for CSMC, demonstrating its efficacy in perfectly recovering the matrix under standard assumptions.
Numerical simulations confirm the theoretical findings, showcasing CSMC’s ability to achieve significant time savings while preserving solution quality compared to nuclear norm minimization algorithms. Moreover, comparisons with alternative methods underscore CSMC’s competitive performance in terms of relative approximation error.
The presented CSMC methods sample columns following a uniform distribution and employ nuclear norm-based matrix completion algorithms. Nevertheless, the framework presented exhibits versatility, accommodating alternative sampling methods and matrix completion techniques. Recent studies have introduced adaptive sampling strategies [36, 34]. The application of such methodologies could effectively address matrices with higher coherence or reduce sample complexity, which we intend to explore in forthcoming research.

### v5

# Randomized Approach to Matrix Completion: Applications in Collaborative Filtering and Image Inpainting

@@ Abstract
We present a novel method for matrix completion, specifically designed for matrices where one dimension significantly exceeds the other. Our Columns Selected Matrix Completion (CSMC) method combines Column Subset Selection and Low-Rank Matrix Completion to efficiently reconstruct incomplete datasets. In each step, CSMC solves a convex optimization problem. We introduce two algorithms to implement CSMC, each tailored to problems of different sizes. A formal analysis is provided, outlining the necessary assumptions and the probability of obtaining a correct solution. To assess the impact of matrix size, rank, and the ratio of missing entries on solution quality and computation time, we conducted experiments on synthetic data. The method was also applied to two real-world problems: recommendation systems and image inpainting. Our results show that CSMC provides solutions of the same quality as state-of-the-art matrix completion algorithms based on convex optimization, while achieving significant reductions in computational runtime.
Keywords Matrix completion  $\cdot$ Column Subset Selection  $\cdot$ Low-rank models  $\cdot$ Image inpainting  $\cdot$ Collaborative filtering  $\cdot$ Nuclear Norm  $\cdot$ Convex Optimization
@@ 1 Introduction
Images, time series, and graphs are just a few examples of data types that can naturally be represented as real-valued matrices. In practice, these matrices often have missing entries due to data collection costs, sensor errors, or privacy considerations. To fully exploit the potential of matrix-based datasets, we need to address the missing data problem. Matrix completion (MC) methods leverage assumptions about underlying patterns in the data, which often manifest as a low-rank property of the matrix being analyzed. This low-rank characteristic reflects inherent correlations or regularities in the data, enabling effective recovery of missing entries.
Matrix completion has found applications in diverse fields such as signal processing, sensor localization, medicine and genomics [1, 2, 3, 4]. Its significance was underscored by the Netflix Prize challenge, which focused on predicting unknown ratings using collaborative filtering based on the preferences of many users [5, 6, 7]. The task involves completing the sparse user-movie rating matrix $\mathbf{M}\in\mathbb{R}^{n_{1}\times n_{2}}$, where $n_{1}$ (the number of users) is significantly larger than $n_{2}$ (the number of movies). This creates an imbalanced matrix in which most entries are missing, as users only rate a small subset of movies.
Singular Value Decomposition (SVD) forms the basis of many MC algorithms. It is commonly used to exploit the high correlation in the data by decomposing the matrix $\mathbf{M}$ as follows:
@@
where $\mathbf{U}\in\mathbb{R}^{n_{1}\times n_{1}}$, $\mathbf{V}\in\mathbb{R}^{n_{2}\times n_{2}}$ are orthogonal, and $\mathbf{\Sigma}\in\mathbb{R}^{n_{1}\times n_{2}}$ is non-negative diagonal matrix of singular values, ordered according to their magnitude. The singular values indicate the strength of the variance along the corresponding directions, which are captured by the columns of $\mathbf{U}$ and $\mathbf{V}$. The rank of $\mathbf{M}$ is determined by the number of non-zero singular values. A common approach to MC is to identify the matrix of lowest rank that best approximates the observed entries. However, directly minimizing the rank is NP-hard [5, 8]. Instead, minimizing the nuclear norm, which is the sum of the singular values of the matrix, is often pursued [5]. Let $\Omega\subseteq\{1,\cdots,n_{1}\}\times\{1,\cdots n_{2}\}$ be the set of indices of the known entries, the convex optimization problem for the exact MC has the following form
@@ 2.3 Our contributions
    We introduce a two-staged matrix completion method, CSMC, dedicated to recovering rectangular matrices with one direction significantly larger than the other ($n_{1}<<n_{2}$).This structure is particularly well-suited for large-scale and asymmetrical data scenarios common in multi-source data integration tasks.
@@
    We validate the CSMC method through numerical experiments on synthetic and real-world datasets, comparing its performance with matrix completion algorithms from the literature, specifically for recommendation systems and image inpainting.
@@ Results
Bridge restored with CSNN-0.7
Figure 8: Image inpainting; NN and CSNN-0.7 algorithms, $\rho=20\%$ known entries.
@@ 5 Conclusion
This paper introduces the Columns Selected Matrix Completion (CSMC) method, which enhances the time efficiency of nuclear norm minimization algorithms for low-rank matrix completion. CSMC employs a two-stage approach: first, it applies a low-rank matrix completion algorithm to a reduced problem size, followed by minimizing the least squares error. Theoretical analysis provides provable guarantees for perfect matrix recovery under standard assumptions.
Numerical simulations show that CSMC significantly reduces computational time while maintaining high accuracy in the reconstructed data. Compared to alternative methods, CSMC demonstrates competitive performance in approximating missing entries. This enhanced time efficiency is particularly valuable for tasks such as recommendation systems, image processing, and other matrix-based applications, where large, incomplete datasets must be effectively recovered. Reducing computational time while minimizing accuracy loss is crucial for enabling efficient processing of incomplete data, which is central to many real-world applications in pattern recognition.
CSMC relies on uniform column sampling and nuclear norm-based algorithms, but its flexible framework allows for the integration of alternative sampling strategies and matrix completion techniques. Future work will explore adaptive sampling approaches, such as those proposed in recent studies [37, 35], to handle high-coherence matrices and reduce sample complexity.